Evaluate integral of a = 0 and b =Ï€\2 sin(t)dt by applying Simpsons rule with four equal intervals


1
Expert's answer
2021-11-12T15:48:21-0500
"\displaystyle\int_{a}^{b}f(t)dt"

"\approx\dfrac{\Delta t}{3}\big(f(t_0)+4f(t_1)+2f(t_2)+4f(t_3)"

"+2f(t_4)+...+4f(t_{n-2})+2f(t_{n-1})+f(t_n)\big)"

"\Delta t=\dfrac{b-a}{n}"

We have that "f(t)=\sin t, a=0, b=\pi/2, n=4."

Therefore


"\Delta t=\dfrac{\pi/2-0}{4}=\pi/8"

Divide the interval "[0, \pi/2]"  into "n=4" subintervals of the length "\u0394t=\u03c0\/8" with the following endpoints: "a=0, \pi/8, \pi/4, 3\pi/8, \pi/2=b."

 Evaluate the function at these endpoints


"f(t_0)=f(0)=\sin(0)=0"


"4f(t_1)=4f(\pi/8)=4\sin(\pi/8)"

"\u22481.530733729460359"


"2f(t_2)=2f(\pi/4)=2\sin(\pi/4)=\sqrt{2}"


"\approx1.414213562373095"


"4f(t_3)=4f(3\pi/8)=4\sin(3\pi/8)"

"\u22483.695518130045147"

"f(t_4)=f(\pi/2)=\sin(\pi/2)=1"

Therefore


"\displaystyle\int_{0}^{\pi/2}\sin t dt"

"\approx(\pi/12)(0+1.530733729460359"

"+1.414213562373095"

"+3.695518130045147+1)"

"\approx1.000134584974194"

"\displaystyle\int_{0}^{\pi/2}\sin t dt\approx1.000134584974194"